OpenStudy (anonymous):

-5 times 7 plus 10 divide by 2 =

OpenStudy (anonymous):

-30

OpenStudy (anonymous):

can you explain how you got that please? thanks

OpenStudy (anonymous):

i used pamdas

OpenStudy (anonymous):

Where are your brackets??? ie: no brackets -5x7+10/2 = -30 if the question was something like (-5x7+3)/2 your answer would be different

OpenStudy (anonymous):

multiply first starting from the left, -5*7=35, then divide 10/2=5, then add -35+5

OpenStudy (anonymous):

it stands for parenthesis, exponent, multiplicaton, division, addition, subtraction...respectively...since u had had no parenthesis i just multiplied first